OpenStudy (anonymous):

A=De^(-0.65t) Find an expression for the amount of drug in the body just after the nth dose has been given.

OpenStudy (callisto):

Do you mind explaining what t is, and how long does it take to apply one does?

OpenStudy (callisto):

*dose

OpenStudy (anonymous):

can u little bit elaborate!

OpenStudy (anonymous):

Hmm well the only stuff I wrote down about it is A=De^-0.65t A=De^-0.65*2t A=De^-0.65*3t A=De^-0.65*4t A=De^-0.65*5t

OpenStudy (anonymous):

Suppose the initial dose, D, of a drug is given at time t=0 and that this dosage is repeated at intervals of every T minutes.

OpenStudy (anonymous):

dy/dt =-ky

OpenStudy (callisto):

what t represent?

OpenStudy (anonymous):

The previous question was Show that the amount of initial does remaining when the second dose is given is De^-kt

OpenStudy (callisto):

*does t represent

OpenStudy (anonymous):

time

OpenStudy (callisto):

so, t/n =T ?

OpenStudy (anonymous):

Hmm I'm not sure, I think I'll just ask a classmate that has already done it.

OpenStudy (callisto):

Hmm, I don't have your exact question, with what the unknowns represent, so it is hard to tell.. :S

OpenStudy (anonymous):

Thanks for trying :D

OpenStudy (anonymous):

wrong group

OpenStudy (callisto):

welcome, sorry that i can't help you :(